A riverboat crosses the river going due north at 8m/s crosses a river going due east at 6m/s. If it takes the boat 18s to cross the river, how far downstream did the boat drift?


1
Expert's answer
2021-02-04T14:39:15-0500

"v_n = 8 m\/s"      "v_e = 6m\/s"           "t = 18s"

"S = vt = \sqrt{v_n^2+v_e^2} t = \sqrt{8^2+6^2} * 18 = 180m"